# An Interpretable Deep-Learning Framework for Predicting Hospital Readmissions From Electronic Health Records

### v1

@@ Abstract
With the increasing availability of patients’ data, modern medicine is shifting towards prospective healthcare. Electronic health records contain a variety of information useful for clinical patient description and can be exploited for the construction of predictive models, given that similar medical histories will likely lead to similar progressions. One example is unplanned hospital readmission prediction, an essential task for reducing hospital costs and improving patient health. Despite predictive models showing very good performances especially with deep-learning models, they are often criticized for the poor interpretability of their results, a fundamental characteristic in the medical field, where incorrect predictions might have serious consequences for the patient health. In this paper we propose a novel, interpretable deep-learning framework for predicting unplanned hospital readmissions, supported by NLP findings on word embeddings and by neural-network models (ConvLSTM) for better handling temporal data. We validate our system on the two predictive tasks of hospital readmission within 30 and 180 days, using real-world data. In addition, we introduce and test a model-dependent technique to make the representation of results easily interpretable by the medical staff. Our solution achieves better performances compared to traditional models based on machine learning, while providing at the same time more interpretable results.
@@ 1 Introduction
Thanks to the availability of a huge amount of data regarding already treated individuals, modern medicine is recently moving towards prospective healthcare, whose goal is to determine the risk for individuals to develop specific diseases, detect and prevent them as soon as possible, and thus provide maximum benefit [1]. Indeed, healthcare data allow to predict health risks for patients with higher accuracy, thus lowering hospitalization costs and improving the effectiveness of medical treatments. Electronic Health Records (EHRs) are among the richest sources of health information. Their usage is steadily growing, with the adoption of EHRs by office-based physicians in the U.S. reaching 88% in 2021 [2]. EHRs are composed of medical episodes, each containing various data of a patient, such as demographic information, diagnoses, clinical notes and more; each episode is associated with a timestamp, thus making it time-dependent. Since diseases tend to form clusters—determined by comorbidity [3]—EHRs can be used for different tasks related to prospective healthcare, such as heart failure prediction [4], suicide risk stratification [5] or general disease prediction of the next diagnoses [6]. However, the information contained in EHRs is also very heterogeneous in nature: the presence of clinical codes (e.g., ICD-9 codes) and datetime objects (e.g., time of discharge) makes EHRs rather challenging [7].
@@
One of the many applications of prospective healthcare is predicting unplanned hospital readmissions; unplanned readmissions consume lots of resources for the hospital, at the same time putting the patient at risk. For example, the U.S. Congress created the Hospital Readmissions Reduction Program [8] to reduce funding for hospitals with excessive readmissions, which is often regarded as a symptom of incorrect treatment or premature discharge [9]. Thanks to the ever-growing amount of clinical data and the constant improvements in the predictive capabilities of deep-learning algorithms, nowadays it is possible to create better decision support systems for predicting hospital readmissions and further reduce patients’ risks. Deep-learning models for healthcare are becoming more and more popular, but at the same time their results are frequently criticized for their lack of interpretability [7]; in the medical field, where incorrect predictions can negatively influence the patients’ health, it is fundamental for predictive models to be sufficiently interpretable for users to understand the rationale behind a decision. Recently, solutions based on Convolutional Neural Networks (CNN) with few layers proved to be accurate in readmission prediction tasks, while at the same time their simple structure leaves room for interpretability analyses [10]. However, CNN are not able to properly assess long-term dependencies, which are particularly important in patients with a long clinical history.
In this paper, we introduce a novel framework for readmission prediction, named ConvLSTM1d, that maintains all the advantages of CNNs while also introducing a memory dimension, thanks to its ConvLSTM architecture based on Long Short-Term Memory (LSTM) to manage long-term dependencies. ConvLSTM1d improves predictions of unplanned hospital readmissions compared to both state-of-the-art methods and traditional models based on machine learning, while providing at the same time interpretable results. Until now, ConvLSTM models have mostly been used in the healthcare field for image segmentation tasks [11, 12], thus making our work the first attempt at leveraging them for tasks of medical prediction. Moreover, we introduce an ad-hoc model-dependent interpretability technique to visualize the prediction results and make them more understandable and easier to read for the medical staff.
@@
This paper is organized as follows: in Section 2 we present works related to our solutions, regarding readmission prediction and interpretability in deep-learning models; in Section 3 we illustrate the architecture of ConvLSTM1d; Section 4 presents experimental and performance results of our model; Section 5 describes the model-dependent technique to make ConvLSTM1d more interpretable; finally, in Section 6 we sum up this work and our contributions.
@@ 2 Related Work
Traditional machine learning techniques to analyze data contained in EHRs (e.g., random forest [13] and logistic regression [14]) often require heavy pre-processing and feature engineering. These steps are time-consuming and require extensive user domain knowledge. In recent years, deep-learning techniques have been increasingly adopted for the analysis of EHRs, thanks to the reduced human contribution required [15]. As pointed out in [6], research in disease predictions can be divided into two areas: specific-purpose progress modeling relates to a specific type of disease, requiring heavy use of domain-specific knowledge; general-purpose progress modeling is instead adaptable for a wide range of diseases. We focus on research works from the second category, since our contribution is also general-purpose.
DeepCare [16] is an LSTM-based framework for outcome prediction from EHRs, used to model long-term dependencies in a variety of healthcare tasks. Outcome prediction is a more general task compared to readmission prediction, but very similar in nature. The authors evaluate DeepCare on three tasks: disease progression, intervention recommendation and readmission prediction (within 3 and 12 months).
Doctor AI [6] is another framework for outcome prediction, based instead on a Recurrent Neural Network (RNN). Predictions of future diagnoses and medications are made by first taking, from the EHR, temporally-ordered codes for past diagnoses and medications, and then converting them into word embeddings to add a semantic dimension. Doctor AI introduces a Top-K recall to mimic the behavior of doctors who would make a list of the K most likely future diagnoses and medications.
The authors of [9] present a framework specifically designed to predict unplanned readmissions within 30 days from ICU discharge. This system uses a large amount of patient data extracted from the MIMIC III database [17]. The proposed model is based on LSTMs, to better handle time series data. In particular, the authors present a preliminary study to analyze and compare a bidirectional LSTM, a LSTM+CNN, and combinations of various features such as word embedding layers. The best performances are obtained by an LSTM+CNN architecture using all the available data. At the time the paper was published, this type of architecture was very new, thus their results do not deviate much from a simple logistic regression, leaving room for improvement.
Finally, Deepr is yet another framework to solve outcome prediction tasks with EHRs [10]. Not only it obtains accurate results, but it also represents a first attempt at creating an interpretable model that detects hidden relationships within the data, showing that interpretable deep learning has great potential in medical informatics. Deepr is based solely on a CNN, unlike other frameworks which use RNNs and LSTMs. Despite a simple architecture, they find space for an embedding layer based on word2vec. Moreover, the authors suggests that a convolutional layer is fundamental to determine the co-occurrence of diagnosis and procedures.
@@
In the works previously presented, pre-trained embeddings are included to provide an additional layer for powerful semantic representations of medical concepts. Models such as word2vec [18, 19] can be trained on specific domains such as healthcare to increase the accuracy of the word embeddings they produce. In this context, the choice of the training data is crucial for obtaining good performance [20]. For instance, [21] introduces a pre-trained embedding model for representing medical concepts, based on word2vec and trained on a dataset containing information of over 4 million patients. The generated vector space includes ICD-9-CM diagnoses and procedures codes, as well laboratory and drug codes from other medical dictionaries.
@@
Finally, the concept of interpretability is widely debated within the scientific community, and it largely depends on the task and context considered [22]. In this paper, we consider as “interpretable” a model for which the reasons that led to a certain decision are made clear and accessible. LIME (Local Interpretable Model-agnostic Explanations) [23] is an example of a framework to make classifiers and regressors more interpretable and reliable. LIME is very flexible and can be used for different types of input, such as images or text. Recent works from the literature, such as [24], studied the interpretability of CNNs applied to text, showing the limits of the traditional approaches that use convolutional filters as n-gram detectors, and a subsequent max-pooling layer to select the most relevant n-grams. As for outcome prediction for EHRs with neural networks, [25] implements an easily interpretable model for predicting unplanned readmissions. Similar to the Deepr framework, this classifier is based on a simple CNN network, which offers relatively average performance but, on the other hand, can be easily interpreted.
@@ 3 ConvLSTM1d
The structural choices that motivated our research, called ConvLSTM1d, are partly ascribable to the limitations highlighted by the authors of the system presented in [10], i.e., the best-performing readmission prediction framework present in literature when we started working on this task: “There are rooms for future work. First, long-term dependencies are simply captured through a max-pooling operation. This is rather simplistic due to a complex dynamic between care processes and disease processes. A better model should pool information that is time-sensitive (e.g., recent events are more important to distant ones).”
Based on these statements, we came up with the following two solutions:
  • Structure medical records in a new way, in order to take care of event timing.
  • Use a new model based on an LSTM architecture that is able to exploit this new structure to “understand" long-term dependencies.
Our framework is composed of two main steps: EHR representation and readmission prediction.
@@ 3.1 EHR Representation
The EHR representation is a crucial part of our methodology. The input of this phase is a data source in which each line uniquely corresponds to a hospital admission, and at the end of this phase the system returns a structured representation of the medical information to be used by the predictive algorithm. To predict a future readmission, the whole patient history plays an important role. To this scope, if an individual has multiple hospital admissions, to construct a useful training set, the patient’s hospitalization at time $T$ must consider all the information about previous timestamps. To include information on time intervals, the system computes the time gap between two consecutive admissions. We define five possible intervals: 0-1 month, 1-3 months, 3-6 months, 6-12 months, more than 12 months.
Each hospitalization is described by a series of diagnoses and procedures, encoded inside the dataset using ICD-9-CM codes, an international classification system for diseases, injuries, surgeries, and diagnostic and therapeutic procedures. [^1] https://www.cdc.gov/nchs/icd/icd9cm.htm The following sequence of codes shows the example of a new admission of a patient with three prior admissions in her/his medical history; please note in italic the codes representing the time gaps between one admission and the next one.
@@
The ICD-9-CM codes of each patient record are randomly shuffled within each admission. The reasons are two: (i) given a single admission, there is no temporal order of the codes within the dataset; (ii) as introduced in [10], shuffling allows to capture many more potential clinical motifs and learn more combinations of codes.
Differently from the readmission prediction frameworks available in the literature, a crucial property of our system is the capability of understanding long-term dependencies, which are of paramount importance to identify patterns in subsequent hospital admissions. To achieve this goal, we now present a novel representation of EHRs. In our framework, data are organized in the following way: each patient hospital admission takes into account, in a cumulative way, also the previous admissions. The latter, however, are divided according to the time order in which they happened. Not only we consider the time gaps between two consecutive admissions, but, contrary to other methods, we also include this piece of information in a separate vector.
@@
For greater clarity, in Fig. 1 we display an example of how we represent admissions extracted from EHRs. Surrounded by a blue dashed line, the vector relative to the codes contained in all the admissions undergone by the patient up to that point in time. The green dashed lines surround the vectors related to the time gaps between two admissions. The first value is always “empty” as it is relative to the patient’s first admission.
Structuring the medical records in this way allows to take into account all the admissions related to the patient with the corresponding codes, without the necessity to make any split of the past data typically using a sliding window. This, along with the deep-learning architecture we use, allows us to correctly assess the contribution of past events on future hospital readmissions.
Last but not least, the adopted EHR representation is flexible and allows to exploit also further medical information in the prediction process: for each new set of data that we want to use, it would be enough to encode them in another vector, taking into account the time order. An example could be the addition of drugs administered in each admission, or data regarding the age or weight of the patient for each past admission. Currently, since the other state-of-the-art methods do not offer this feature, we limit ourselves to using only procedural and diagnostic data with time gaps.

### v2

@@ Abstract
With the increasing availability of patient data, modern medicine is shifting towards prospective healthcare. Electronic health records offer a variety of information useful for clinical patient characterization and the development of predictive models, given that similar medical histories often lead to analogous health progressions. One application is the prediction of unplanned hospital readmissions, an essential task for reducing healthcare costs and improving patient outcomes. While predictive models demonstrate strong performances especially with deep learning approaches, they are often criticized for their lack of interpretability, a critical requirement in the medical domain where incorrect predictions may have severe consequences for patient safety. In this paper, we propose a novel and interpretable deep learning framework for predicting unplanned hospital readmissions, supported by NLP findings on word embeddings and by ConvLSTM neural networks for better handling temporal data. We validate the framework on two predictive tasks for hospital readmission within 30 and 180 days, using real-world data. Additionally, we introduce and evaluate a model-dependent technique designed to enhance result interpretability for medical professionals. Our solution outperforms traditional machine learning models in prediction accuracy while simultaneously providing more interpretable results.
@@ 1 Introduction
With the increasing availability of extensive data on previously treated patients, modern medicine is progressively shifting towards prospective healthcare, which aims to determine the risk for individuals for specific diseases, enable early detection and prevention, and maximize patient outcomes [1]. In fact, the richness of healthcare data can be leveraged to predict health risks for incoming patients with high accuracy, bringing critical benefits to both healthcare structures and patients, by lowering hospitalization costs and improving the effectiveness of medical treatments. Electronic Health Records (EHRs) are among the richest sources of health information. This is evident in their steadily growing application, as reflected by the adoption of electronic health records EHRs by office-based physicians in the U.S. reaching 88% in 2021 [2]. EHRs are composed of medical episodes, each containing a series of data about a patient, such as demographic information, biochemical profiles, past and present medical diagnoses, clinical notes, and more. Each patient episode is associated with a timestamp, thus making the information contained in EHRs time-dependent. Diseases often occur in interconnected patterns, forming clusters influenced by comorbidity occurring due to shared risk factors, biological mechanisms, or lifestyle influences [3]. This phenomenon enables EHRs to support a wide range of tasks in prospective healthcare, such as predicting heart failure [4], suicide risk [5], mortality [6], or future diagnoses [7]. However, information contained in EHRs is also very heterogeneous in nature: the presence of clinical codes (e.g., ICD-9 codes) and datetime objects (e.g., time of discharge) makes EHRs rather challenging to use effectively [8].
@@
One of the many applications of prospective healthcare is the prediction of unplanned hospital readmissions. Unplanned readmissions have a significant impact on healthcare systems by consuming many resources, while simultaneously exposing patients to additional health risks. To reduce and discourage unplanned hospital readmissions, governing bodies go so far as to implement regulatory measures. For instance, the U.S. Congress created the Hospital Readmissions Reduction Program [9] to reduce funding for hospitals with excessive readmissions, which is often regarded as a symptom of incorrect treatment or premature discharge [10]. With the ever-increasing volume of clinical data and the continuous improvements in the predictive capabilities of deep learning algorithms, it is now possible to develop more effective decision support systems for predicting hospital readmissions and further reducing patient risks. However, a major concern with the introduction of deep learning models in healthcare is their lack of transparency and interpretability, leading to frequent criticism of their results [8, 11]. Particularly in the medical field, where incorrect predictions can severely affect patient health, it is crucial for predictive models to be sufficiently transparent and interpretable, allowing practitioners to understand the rationale behind decisions and to avoid both ethical and practical risks [12, 13, 14].
Recently, solutions based on Convolutional Neural Networks (CNN) with few layers proved to be accurate in readmission prediction tasks, while their simple architecture also offers significant interpretability benefits [15]. While more complex architectures, such as Transformers [16], achieve state-of-the-art results in many applications, their computational cost can make model training and inference challenging.111 Transformer architectures are computationally intensive due to the quadratic complexity of their self-attention mechanism, which scales as $O(n^{2}\cdot d)$ with respect to the input sequence length $n$ and embedding dimension $d$. While efficient attention variants have been proposed to reduce this complexity [17, 18, 19, 20], their robustness is debated [21]. Additionally, the large number of parameters in Transformer-based models contributes significantly to both memory usage and training/inference cost [22, 23]. Although they are highly parallelizable and there exist techniques to improve memory usage [24, 25], these advantages may not fully compensate for their computational demands in resource-constrained domains such as healthcare. Additionally, their interpretability relies on the often debated explainability of attention mechanisms [26, 27]. Simpler models like CNNs reduces computational complexity while also providing transparent and interpretable results when combined with saliency methods [28], which is particularly important in healthcare applications. However, a critical limitation of CNNs for predicting hospital readmission is the inability to properly assess long-term dependencies, which are particularly important for patients with a long clinical history [29]. To overcome this limitation, in this paper we introduce a novel framework for readmission prediction, named ConvLSTM1d. ConvLSTM1d retains all the advantages of CNNs, such as low complexity and enhanced interpretability, while also incorporating a memory component through its ConvLSTM architecture, which leverages Long Short-Term Memory (LSTM) units to effectively capture long-term dependencies. Our solution improves predictions of unplanned hospital readmissions compared to traditional machine learning models, while simultaneously boosting interpretability with respect to more complex deep learning architectures. Until now, ConvLSTM models have mostly been used in the healthcare field for image segmentation tasks [30, 31]. Few notable works pioneered the use of ConvLSTM in conjunction with EHRs, but without focusing on the interpretability of the results [32, 33]. In this paper, we also introduce and evaluate a model-dependent interpretability approach to visualize predictions, making them more understandable and accessible to medical practitioners.
@@
This paper is organized as follows: in Section 2 we present works related to our solutions, regarding readmission prediction and interpretability in deep learning models; in Section 3 we illustrate the architecture of ConvLSTM1d; Section 4 describes the model-dependent technique to make ConvLSTM1d more interpretable; Section 5 presents the experimental setting and the performance results of our model; finally, in Section 6 we sum up this work and our contributions.
@@ 2 Related Work
Traditional machine learning techniques to analyze data contained in EHRs (e.g., random forest [34] and logistic regression [35]) often require heavy pre-processing and feature engineering. These steps are time-consuming and require extensive user domain knowledge. In recent years, deep learning techniques have been increasingly adopted for the analysis of EHRs, with the reduced human contribution required [8]. As pointed out by [7], research in disease predictions can be divided into two areas: specific-purpose progress modeling relates to a specific type of disease, requiring heavy use of domain-specific knowledge; general-purpose progress modeling is instead adaptable for a wide range of diseases. We focus on research works from the second category, since our contribution is also general-purpose.
DeepCare [36] is an LSTM-based framework for outcome prediction from EHRs, used to model long-term dependencies in a variety of healthcare tasks. Outcome prediction is a more general task compared to readmission prediction, but very similar in nature. The authors evaluate DeepCare on three tasks: disease progression, intervention recommendation and readmission prediction (within 3 and 12 months). Doctor AI [7] is another framework for outcome prediction, based instead on a Recurrent Neural Network (RNN). Predictions of future diagnoses and medications are made by first taking, from the EHR, temporally-ordered codes for past diagnoses and medications, and then converting them into word embeddings to add a semantic dimension. Doctor AI introduces a Top-K recall to mimic the behavior of doctors who would make a list of the K most likely future diagnoses and medications. [10] present a framework specifically designed to predict unplanned readmissions within 30 days from ICU discharge. This system uses a large amount of patient data extracted from the MIMIC III database [37]. The proposed model is based on LSTMs, to better handle time series data. In particular, the authors present a preliminary study to analyze and compare a bidirectional LSTM, a LSTM+CNN, and combinations of various features such as word embedding layers. The best performances are obtained by an LSTM+CNN architecture using all the available data. At the time the paper was published, this type of architecture was very new, thus their results do not deviate much from a simple logistic regression, leaving room for improvement. Finally, Deepr is yet another framework to solve outcome prediction tasks with EHRs [15]. Not only it obtains accurate results, but it also represents a first attempt at creating an interpretable model that detects hidden relationships within the data, showing that interpretable deep learning has great potential in medical informatics. Deepr is based solely on a CNN, unlike other frameworks which use RNNs and LSTMs. Despite a simple architecture, they find space for an embedding layer based on Word2Vec. Moreover, the authors suggests that a convolutional layer is fundamental to determine the co-occurrence of diagnosis and procedures.
@@
In the works previously presented, pre-trained embeddings are included to provide an additional layer for powerful semantic representations of medical concepts. Models such as Word2Vec [38, 39] can be trained on specific domains such as healthcare with little effort, to increase the accuracy of the word embeddings they produce. In this context, the choice of the training data is crucial for obtaining good performance [40]. For instance, [41] introduce a pre-trained embedding model for representing medical concepts, based on Word2Vec and trained on a dataset containing information of over 4 million patients. The generated vector space includes ICD-9-CM diagnoses and procedures codes, as well laboratory and drug codes from other medical dictionaries. More recently, Transformer-based models such as BERT, which produce contextualized embeddings, have been included in risk prediction of patient hospital readmission [29, 42, 43]. Unlike traditional methods that rely on static embeddings, which represent words in a fixed manner, BERT models generate embeddings that are dynamic and context-dependent, potentially offering a more flexible approach, at the expense of increasing complexity in terms of number of parameters, layers, and size of the training data.
@@
Finally, the concept of interpretability is widely debated within the scientific community, and it largely depends on the task and context considered [44]. In this paper, we consider as “interpretable” a model for which the reasons that led to a certain decision are made clear and accessible. Recent literature reviews on the use of EHR for predicting hospital readmission emphasize the critical need to address interpretability in medicine [45]. The evolution of this field necessitates the development and implementation of interpretable machine learning methods to enhance clinical usefulness. LIME (Local Interpretable Model-agnostic Explanations) [46] is an example of a framework to make classifiers and regression models more interpretable and reliable. LIME is very flexible and can be used for different types of input, such as images or text. Recent works from the literature, such as [47], studied the interpretability of CNNs applied to text, showing the limits of the traditional approaches that use convolutional filters as n-gram detectors, and a subsequent max-pooling layer to select the most relevant n-grams. As for outcome prediction for EHRs with neural networks, the solution described by [48] implements an easily interpretable model for predicting unplanned readmissions. Similar to the Deepr framework, this classifier is built on a simple CNN architecture, which delivers relatively average performance but offers the significant advantage of being easily interpretable.
Recently, [42] introduced a model of the BERT family for predicting patient admissions based on emergency department triage notes. The authors leverage LIME to provide interpretable predictions through visualization of influential words in the notes, showing that combining NLP with explainable AI can assist in admission decisions. BEHRT [29] introduces a novel Transformer-based methodology for EHR data by combining embeddings—disease codes, age, positional info, and visit segments to pre-train the model on a masked language task. BEHRT captures complex temporal and contextual dependencies, offering both interpretable results through attention visualization and higher predictive accuracy compared to traditional methods like RNNs and CNNs. ExBEHRT [43] extends the feature space of BEHRT to several multi-modal records, such as demographics, vital signs, and medications. At the same time, interpretability of Transformer-based models is frequently debated in recent literature [27]. Whether attention mechanisms truly provide meaningful explanations in terms of faithfulness and plausibility is still an open question [26]. Moreover, methods such as LIME can already boost interpretability of deep learning models, suggesting that attention visualization may not be needed [28].
Traditional machine learning methods remain a subject of study for hospital readmission prediction tasks, due to their ease of implementation, lower resource costs, and the favorable trade-off between transparency and performance. [49] recently introduced two traditional models for readmission prediction, based on logistic regression and gradient boosting machine, that achieve good performance by integrating manually engineered features with Word2Vec features derived from administrative health data.
@@ 3 ConvLSTM1d
The structural choices that motivated our research output, namely ConvLSTM1d, are partly ascribable to the limitations highlighted by [15] and their system, one of the best-performing readmission prediction frameworks present in the literature at the beginning of our research. In their solution, long-term dependencies are captured through a max-pooling operation, which is a simplistic approach given the complex dynamics between care processes and disease processes. A more effective model should prioritize time-sensitive information, where recent events are given more weight than distant ones. Based on these statements, we came up with the following two solutions: (i) structure medical records in a new way, in order to take care of event timing, and (ii) use a new model based on an LSTM architecture that is able to exploit this new structure to “understand" long-term dependencies.
Our framework consists of two main steps: EHR representation and readmission prediction.
@@ 3.1 EHR Representation
The EHR representation is a crucial part of our methodology. The input of this phase is a data source in which each line uniquely corresponds to a hospital admission, and at the end of this phase the system returns a structured representation of the medical information to be used by the predictive algorithm. To predict a future readmission, the whole patient history plays an important role. In this context, if an individual has multiple hospital admissions, to construct a useful training set, the patient’s hospitalization at time $T$ must consider all the information about previous timestamps. To include information on time intervals, the system computes the time gap between two consecutive admissions. We define five possible intervals: 0-1 month, 1-3 months, 3-6 months, 6-12 months, and more than 12 months.
Each hospitalization is described by a series of diagnoses and procedures, encoded within the dataset using ICD-9-CM codes, an international classification system for diseases, injuries, surgeries, and diagnostic and therapeutic procedures. [^2] https://www.cdc.gov/nchs/icd/icd9cm.htm The following sequence of codes shows the example of a new admission of a patient with three prior admissions in her/his medical history; please note in italic the codes representing the time gaps between one admission and the next.
@@
The ICD-9-CM codes for each patient record are randomly shuffled within each admission. The reasons are twofold: (i) given a single admission, there is no temporal order of the codes within the dataset; (ii) as introduced by [15], shuffling records allows to capture many more potential clinical motifs and learn more combinations of codes.
Unlike a number of readmission prediction frameworks available in the literature, a crucial property of our system is the ability to understand long-term dependencies, which are of paramount importance to identify patterns in subsequent hospital admissions [29]. To achieve this goal, we introduce a novel representation of EHRs. In our framework, data are organized in the following way: each patient hospital admission takes into account, in a cumulative way, also the previous admissions. The latter, however, are divided according to the time order in which they happened. Not only we consider the time gaps between two consecutive admissions, but, contrary to other methods, we also include this piece of information in a separate vector.
@@
For greater clarity, in Fig. 1 we display an example of how we represent admissions extracted from EHRs. The vector representing the codes contained in all admissions undergone by the patient up to that point in time is surrounded by a blue dashed line, while the green dashed lines highlight the vectors corresponding to the time gaps between consecutive admissions. The first value is always “empty” as it is relative to the patient’s first admission.
Structuring the medical records in this way allows to take into account all the admissions related to the patient with the corresponding codes, without the necessity to make any split of the past data, typically by employing a sliding window. This, along with the deep learning architecture adopted, allows to correctly assess the contribution of past events on future hospital readmissions. Finally, the adopted EHR representation offer a high degree of flexibility, by enabling the integration of additional medical information into the prediction process. To incorporate new data, it is sufficient to encode it into a separate vector, ensuring the temporal order is preserved. For instance, this could be the inclusion of drugs administered during each admission or patient-specific data such as age or weight from previous admissions. Currently, since the other state-of-the-art methods do not offer this feature, we limit ourselves to using only procedural and diagnostic data with time gaps.